Parth (parthkohli):

@ganeshie8 @mathslover

Parth (parthkohli):

A family of lines is given by \((1 + 2\lambda) x + (1 - \lambda) y + \lambda = 0 \), where \(\lambda\) is the parameter. The line belonging to this family at the maximum distance from the point \((1,4)\) is?

OpenStudy (anonymous):

what is the question anonymous person?

Parth (parthkohli):

Here is what I did:\[(x + y) + \lambda (2x - y + 1) = 0\]So the point of concurrency is \((1/3, -1/3)\). Now I wrote the line as\[y + 1/3 = m(x - 1/3)\]

Parth (parthkohli):

Then I calculated the distance in terms of \(m\). But I have no idea how one maximizes or minimizes that.

Parth (parthkohli):

Oh, whoops. The point of concurrency is \((-1/3,1/3)\)

Parth (parthkohli):

\[y + 1/3 = m(x - 1/3)\]\[y - mx + \frac{m+1}{3} = 0\]

Parth (parthkohli):

\[d = \dfrac{|4 - m + (m+1)/3|}{\sqrt{m^2 + 1}}\]

Parth (parthkohli):

...aaaand I'm stuck. Any other ways to do this? Maybe keeping things in terms of lambda?

Parth (parthkohli):

where's mathslover when I need him

Parth (parthkohli):

considering different intervals and differentiating may help...

ganeshie8 (ganeshie8):

basically you're given a family of lines all passing through a fixed point and asked to find the maximum distance from a point your method works perfectly!

Parth (parthkohli):

Just that I don't know how to maximise/minimise. :|

Parth (parthkohli):

OK, let's try it interval-by-interval.

Parth (parthkohli):

\[4 - m + (m+1)/3 \ge 0 \]\[12 - 3m + m + 1 \ge 0\]\[2m \le 13 \]\[m \le 13/2\]So that's the case one, and our expression is given by\[\dfrac{\frac{m+1}{3} + 4 - m}{\sqrt{m^2 + 1}}\]

ganeshie8 (ganeshie8):

i think we can try maximizing square of distance to avoid the cases

Parth (parthkohli):

ah, of course. why didn't I think of that?

Parth (parthkohli):

\[\dfrac{m+1}{3} - m + 4 = \dfrac{-2m}{3} + \dfrac{13}{3}\]Squaring,\[\dfrac{4}{9}m^2 - \dfrac{52}{9} m + \frac{169}{9}\]Because we're maximizing, we can remove the fractions and stuff.\[f(m)= \frac{4m^2 - 52m + 169}{m^2 + 1}\]is what we need to maximize.

Parth (parthkohli):

So this is a good-old rational expression... ... which, unfortunately, I'm too lazy to work with. WolframAlpha overlords!

Parth (parthkohli):

m = -2/13. Yup.

mathslover (mathslover):

Hey Parth, sorry, was busy in some other work. Didn't notice that. I am looking at your work now. Give me 2 minutes, sorry for the delay yaar!

Parth (parthkohli):

\[y +1/3 = -2/13(x -1/3)\]

Parth (parthkohli):

brb, battery low

ganeshie8 (ganeshie8):

i wonder how you get \(\lambda \) out of \(m\) hmm

Parth (parthkohli):

comparing coefficients?

Parth (parthkohli):

\[13y + 13/3 = -2x + 2/3\]\[2x + 13y + 11/3 = 0\]\[6x + 39y + 11 = 0\]which is of course the incorrect answer. :)

mathslover (mathslover):